# Off-Policy Evaluation of Slate Bandit Policies via Optimizing Abstraction

### 03 Feb 2024 v1

@@ 6. Conclusion and Future Work
This paper studied OPE of slate contextual bandits where existing estimators encounter significant challenges in terms of bias and variance due to combinatorial action spaces and restrictive assumptions on the reward function. To overcome these limitations, we proposed LIPS, a novel estimator that is built on slate abstraction to substantially reduce variance. Our analysis demonstrates that LIPS can be unbiased given a sufficient abstraction, and provides a substantial reduction in variance. We also observed that the advantages of LIPS may be maximized when we intentionally use an insufficient abstraction. Based on this analysis, we presented a method for optimizing slate abstraction to minimize the bias and variance of LIPS, which differentiates our work from MIPS (Saito and Joachims, 2022b) and OffCEM (Saito et al., 2023). Our experiments illustrate that LIPS enables considerably more accurate OPE for slate actions than existing estimators, including IPS, PI, MIPS, and their DR variants.
As future work, it would be practically valuable to incorporate control variates into LIPS to further improve its variance, which is considered a non-trivial extension since it would change the way we should optimize slate abstractions. Lastly, it would be intriguing to explore the application of recent advances in diffusion models (Yang and Wang, 2023; Abstreiter et al., 2021) to optimize slate abstractions more effectively.
@@ A.5. Additional ablation experiments with varying values of $\beta$
Figure 10. Comparison of the estimators’ MSE (normalized by the true value $V(\pi)$) with varying slate sizes ($L$) and with reward functions (1) - (3) on the Delicious dataset.
Figure 11. Comparing the LIPS’ MSE (normalized by the true value $V(\pi)$) with those of DR estimators under varying slate sizes ($L$) and with reward functions (1) - (3) on the Delicious dataset.
Figure 12. Comparison of the estimators’ MSE (normalized by the true value $V(\pi)$) with varying data sizes ($n$) and with reward functions (1) - (3) on the Delicious dataset.
Figure 13. Comparing the LIPS’ MSE (normalized by the true value $V(\pi)$) with those of DR estimators under varying data sizes ($n$) and with reward functions (1) - (3) on the Delicious dataset.

### 17 Feb 2024 v2

@@ 6. Conclusion and Future Work
This paper studied OPE of slate contextual bandits where existing estimators encounter significant challenges in terms of bias and variance due to combinatorial action spaces and restrictive assumptions on the reward function. To overcome these limitations, we proposed LIPS, a novel estimator that is built on slate abstraction to substantially reduce variance. Our analysis demonstrates that LIPS can be unbiased given a sufficient abstraction, and provides a substantial reduction in variance. We also observed that the advantages of LIPS may be maximized when we intentionally use an insufficient abstraction. Based on this analysis, we presented a method for optimizing slate abstraction to minimize the MSE of LIPS, which differentiates our work from MIPS (Saito and Joachims, 2022b) and OffCEM (Saito et al., 2023).
As future work, it would be practically valuable to incorporate control variates into LIPS to further improve its variance, which is considered a non-trivial extension since it would change the way we should optimize slate abstractions. It would also be valuable to perform real-world experiments with slate bandit data from live enviornments once such a public dataset becomes available.
@@ A.5. Additional ablation experiments with varying values of $\beta$
Figure 10. Comparison of the estimators’ MSE (normalized by the true value $V(\pi)$) with varying slate sizes ($L$) and with reward functions (1) - (3) on the Delicious dataset. The y-axis is on a log scale.
Figure 11. Comparing the LIPS’ MSE (normalized by the true value $V(\pi)$) with those of DR estimators under varying slate sizes ($L$) and with reward functions (1) - (3) on the Delicious dataset. The y-axis is on a log scale.
Figure 12. Comparison of the estimators’ MSE (normalized by the true value $V(\pi)$) with varying data sizes ($n$) and with reward functions (1) - (3) on the Delicious dataset. The y-axis is on a log scale.
Figure 13. Comparing the LIPS’ MSE (normalized by the true value $V(\pi)$) with those of DR estimators under varying data sizes ($n$) and with reward functions (1) - (3) on the Delicious dataset. The y-axis is on a log scale.